OpenStudy (anonymous):

A triangle has the following measurements. What is a possible length for the third side? GH = 19, HJ = 8, JG = ?

OpenStudy (asnaseer):

 In the triangle, if angel "x" is zero, then JG=19-8=11 if angel "x" is 180, then JG=19+8=27 so JG must be between 11 and 27